# Private Evolution Converges

### v1

@@ D.2 Upper bounds
[…]
@@
$\sup_{f\in\cal F}\frac{1}{n}\bigg{|}\sum_{i\in[n]}x_{i}[\pi_{k,f}(z_{i})-\pi_{k-1,f}(z_{i})]\bigg{|}\leq\sup_{f_{k}\in T_{k},f_{k-1}\in T_{k-1}}\frac{1}{n}\bigg{|}\sum_{i\in[n]}x_{i}[f_{k}(z_{i})-f_{k-1}(z_{i})]\bigg{|}.$
We will bound the last term above. We have that for any $f_{k}\in T_{k},f_{k-1}\in T_{k-1},i\in[n]$
$\frac{1}{n}|f_{k}(z_{i})-f_{k-1}(z_{i})|\leq\frac{1}{n}|f_{k}(z_{i})-f(z_{i})|+\frac{1}{n}|f(z_{i})-f_{k-1}(z_{i})|\leq\frac{\varepsilon_{k}+\varepsilon_{k-1}}{n}\leq\frac{3\varepsilon_{k}}{n}.$
Hence, $\frac{1}{n}\sum_{i\in[n]}x_{i}[f_{k}(z_{i})-f_{k-1}(z_{i})]$ is a $(3\varepsilon_{k}/\sqrt{n})$-subGaussian random variable. We conclude that
$\mathbb{E}\left[\max_{f_{k}\in T_{k},f_{k-1}\in T_{k-1}}\frac{1}{n}\bigg{|}\sum_{i\in[n]}x_{i}[(f_{k}(z_{i})-f_{k-1}(z_{i})]\bigg{|}\right]\leq C\frac{\varepsilon_{k}\sqrt{\log(\mathcal{N}(\mathcal{F},\|\cdot\|_{\infty};\varepsilon_{k}))}}{\sqrt{n}},$
@@
[…]
@@
From [22] (Lemma 4.2), we know that
@@
where $\mathcal{F}_{L}$ is the class of $L$-Lipschitz functions mapping $\Omega$ to $[-1,1]$. It is easy to see that $D\mathcal{F}_{1/D}={\cal F}_{BL}$, since ${\cal F}_{BL}$ contains all the $1$-Lipschitz functions mapping $\Omega$ to $[-D,D]$. Then, for any $\beta\in[0,D]$, $\log(\mathcal{N}({\cal F}_{BL},\|\cdot\|_{\infty};\beta))$ can be bounded as follows
@@
where the last inequality uses the fact that $\beta\in[0,D]$. Replacing this expression in the bound given by Lemma D.4 we get
[…]
@@
[…]
@@
[…]
@@
[…]
@@
The convergence of empirical measures in $1$-Wasserstein is a topic of interest on its own, that has been studied in the past. We obtain the following corollary from [29]. We note the results in there hold more generally for Banach Spaces.
Lemma D.5 (Corollary of Theorem 3.1 in [29]).
@@
Appendix E Simulations with histogram from Section 4.2
We follow the same simulation setup from Section 6. Namely, we consider the sample space to be the unit $\ell_{2}$ ball in $\mathbb{R}^{2}$. We use privacy parameters $\varepsilon=1,\delta=10^{-4}$. We set $T=2\log(n\varepsilon)$, and the noise $\sigma$ is computed with the analytic Gaussian mechanism ([4], Theorem 8). The rest of the parameters are set as indicated in Theorem 4.1. We run Algorithm 2, with the only difference that we post-process noisy histograms truncating at $H=0$ and then re-normalize. In Figure 4, we refer to this algorithm as ‘PE’.
@@
Figure 4: Comparison of the performance of PE with different histograms. Both algorithms, ‘PE’ and ‘PE with Laplace noise+thresholding’ are described in Section E. Left: the sensitive dataset is distributed uniformly in an $\ell_{2}$ ball of radious $0.02$ around the origin (heavily clustered), leading ‘PE with Laplace noise+thresholding’ to preserve the votes signal in the noisy NN histograms, outperforming ‘PE’. Right: the sensitive dataset is distributed uniformly in an $\ell_{2}$ ball of radious $0.5$ around the origin (not clustered). The votes signal in the noisy NN histograms is lost by ‘PE with Laplace noise+thresholding’ during the thresholding step. Hence, it performs poorly. Results are averaged over 100 runs
Appendix F Extension to Banach Spaces
@@
Theorem F.1 (Convergence of PE).
Let $(\Omega,\rho)$ be a (sample) Banach space. Suppose $\Omega$ is compact. Let $01>\gamma>$ 0, $v\in\mathbb{N}$ and $\alpha>0$. $S_{T}$ be the output of Algorithm 2 run on input $S\in\Omega^{n},T\in\mathbb{N},n_{s}\in\mathbb{N},\sigma>0,\rho(\cdot,\cdot)$. If $\operatorname{Variation\_API}(\cdot)$ is an $(\gamma,v,\alpha)$-API for $S$, then
@@
The upper bound on the Gaussian complexity that we provided in Lemma D.4 can be used to find $\hat{G}_{vn_{s}}({\cal F}_{BL})$, since it works for general metric spaces, and the upper bound $\hat{W}_{1}(n_{s}))$, can be found in [29].
The proof of Theorem F.1 follows similarly to Theorem 4.1. We provide it below for completeness.
Proof of Theorem F.1.
Note that for any $t=0,...,T-1$
@@
where the inequality follows from $D_{BL}(\hat{\mu}_{t+1},\tilde{\mu}_{t+1})\geq D_{BL}(\tilde{\mu}_{t+1},\mu^{\prime}_{t+1})$ by definition of $\mu^{\prime}_{t+1}$. Denote by $\mathbb{E}_{t}[\cdot]$ the expectation when conditioning on the randomness up to iteration $t$ of PE. Since $\operatorname{Variation\_API}(\cdot)$ is an $(\gamma,v,\alpha)$-API, then Lemma D.2 gives
@@
Further, since $|V_{t}|=vn_{s}$, Lemma D.3 and the definition of $\hat{G}_{n_{s}}({\cal F}_{BL})$ imply
@@
Appendix G Missing proofs
@@
  • If $\hat{\mu}[i]>0$, then $\tilde{\mu}[i]=\hat{\mu}[i]+L_{i}$, where $L_{i}\sim\text{Lap}(2/n\varepsilon)$. If $\tilde{\mu}[i]<2\log(1/\delta)/(n\varepsilon)+1/n$ (i.e the coordinate is small), then it is truncated: $\tilde{\mu}[i]=0$
@@
  • We continue by bounding $D_{BL}(\nu,\tilde{\mu})$.
    $\mathbb{E}[D_{BL}(\nu,\tilde{\mu})]$ $=\mathbb{E}\left[\sup_{f\in{\cal F}_{BL}}\sum_{i\in\tilde{I}}f(V[i])(\hat{\mu}[i]+L_{i}-(\hat{\mu}[i]+L_{i})\mathbbm{1}_{(\hat{\mu}[i]+L_{i}\geq H)})\right]$
@@
For simplicity denote the packing number $M(\Omega,\rho;2\eta)$ by $M$ in this proof. Consider a $2\eta$-packing of $\Omega$, $\{z_{1},...,z_{M}\}$. Let $S=\{z_{1}\}^{n}$. For $k\in\mathbb{Z}_{+}$, define
@@
It is easy to see that if $S=_{\eta}\mathcal{A}(S)\iff\mathcal{A}(S)\in B(1)$, and if $\mathcal{A}(S)\in B(k)$ for some $k\geq 2$, then $S$ and $\mathcal{A}(S)$ can not be $\eta$-close. Since $\mathbb{P}_{\mathcal{A}}\bigg{[}S=_{\eta}\mathcal{A}(S)\bigg{]}\geq 1-\tau$, then $\mathbb{P}_{\mathcal{A}}\bigg{[}\mathcal{A}(S)\in\cup_{k\in\mathbb{Z}\cap[2,M]}B(k)\bigg{]}\leq\tau$. Furthermore, $B(i)\cap B(j)=\emptyset$ for all $i\neq j\in\mathbb{Z}\cap[2,M]$. Hence, there exists a set $B(k^{*})$ such with $\mathbb{P}_{\mathcal{A}}\bigg{[}\mathcal{A}(S)\in B(k^{*})\bigg{]}\leq\tau/(M-1)$. Finally, construct $S^{\prime}=\{z_{1}\}^{n-1}\cup\{z_{k^{*}}\}$. Note $S$ and $S^{\prime}$ are neighboring datasets. Then, by the inequalities that we have stated and $\varepsilon$-DP, it follows that
$1-\tau\leq\mathbb{P}_{\mathcal{A}}\bigg{[}S^{\prime}=_{\eta}\mathcal{A}(S^{\prime})\bigg{]}=\mathbb{P}_{\mathcal{A}}\bigg{[}\mathcal{A}(S^{\prime})\in B(k^{*})\bigg{]}\leq e^{\varepsilon}\mathbb{P}_{\mathcal{A}}\bigg{[}\mathcal{A}(S)\in B(k^{*})\bigg{]}\leq\frac{e^{\varepsilon}\tau}{M-1}.$
Solving for $\varepsilon$, we get $\varepsilon\geq\log((M-1)(1-\tau)/\tau)$. ∎
@@
Appendix H Details on the Private Signed Measure Mechanism
The Private Signed Measure Mechanism is an algorithm for differentially private synthetic data generation presented in [24]. It is designed to work under pure differential privacy. For completeness, we provide an extension of it that works under approximate DP. The algorithm and its proof closely resemble the one from [24], with the only difference being that we need to deal with Gaussian noises instead of Laplace in order to achieve approximate DP. As we mentioned in Section 5, PE can be seen as a sequential version of PSMM, and the gaussian mechanisms has better composition guarantees than the Laplace mechanism. Hence, even though changing Gaussian noise by Laplace in PSMM does not make a big difference, it does make a more notorious difference for PE when seen as a sequential version of PSMM.
@@ Appendix H Details on the Private Signed Measure Mechanism
2:  Let $\tilde{\mu}$ be a signed measure such that for $i\in[m]$, $\tilde{\mu}(\{\omega_{i}\})=\tilde{n}_{i}/n$ for an arbitrary $\omega_{i}\in\Omega_{i}$ and $\tilde{\mu}(\Omega_{i}\backslash\{\omega_{i}\})=0$.
@@
Theorem H.1 (Guarantees of PSMM under approximate DP).
@@
Remark H.1.
When running PSMM, we can select $n_{s}$ arbitrarily large, by the post-processing property of DP. Hence, we can safely assume that the error term $D\left(\frac{\sqrt{\log(1/\delta)}}{n\varepsilon}\right)^{1/\max\{2,d\}}$ dominates in the utility bound presented in Theorem H.1. Note that, up to logarithmic factors, this bound improves over the bound that we provide for PE in Theorem 4.1 by a factor of $d$. However, it requires $\{\Omega_{i}\}_{i\in[m]}$ to be such that $\max_{i\in[m]}\operatorname{diam}(\Omega_{i})=O(Dm^{-1/\max\{d,2\}})$. Constructing such partition in practice is equivalent to finding an $O(Dm^{-1/\max\{d,2\}})$-net of points $\{\omega_{1},....,\omega_{m}\}$ such that the Voronoi partition induced by them is $\{\Omega_{i}\}_{i\in[m]}$. We explained in Section 5 why such net can be difficult to construct in practice.
@@
The convergence proof is similar to the utility analysis of one step of PE. Let $\bar{\mu}$ be a probability measure over $\Omega$ such that $\bar{\mu}(\{\omega_{i}\})=n_{i}/n$, where $n_{i}=|S\ \cap\ \Omega_{i}|$, and $\bar{\mu}(\Omega\backslash\{\omega_{1},...,\omega_{m}\})=0$. Let the measures $\tilde{\mu},\hat{\mu}$ and $\mu$ be as given by Algorithm 5. Note that
@@
$\leq 2\left(\max_{i\in[m]}\operatorname{diam}(\Omega_{i})+D_{BL}(\bar{\mu},\tilde{\mu})\right),$
where the second inequality follows from the fact that $D_{BL}(\bar{\mu},\tilde{\mu})\geq D_{BL}(\tilde{\mu},\hat{\mu})$, which is a consequence of $\hat{\mu}$ being the closest probability measure over $\{\omega_{1},...,\omega_{m}\}$ to $\tilde{\mu}$ in $D_{BL}$, and in the last inequality we used $\max\{W_{1}(\mu_{S},\bar{\mu}),W_{1}(\hat{\mu},\mu)\}\leq\max_{i\in[m]}\operatorname{diam}(\Omega_{i})$ (these inequalities are trivial since the pairs of measures $\mu_{S},\bar{\mu}$ and $\hat{\mu},\mu$ assign the same amount of probability mass into each region $\Omega_{i}$). It remains to bound $D_{BL}(\bar{\mu},\hat{\mu})$. The following equality
@@
For the second claim, recall from Corollary D.1 that
@@
Finally, since $S^{\prime}\sim\mu^{n_{s}}$, Lemma D.5 gives the following upper bound on the convergence of the empirical measure in $W_{1}$
@@
Appendix I Impact Statement

### v2

@@ E.2 Upper bounds
$G_{n}(\mathcal{F})\leq C\inf_{\alpha\geq 0}\left[\alpha+\int_{\alpha}^{D}\sqrt{\log(\mathcal{N}(\mathcal{F},\|\cdot\|_{\infty};\beta))}d\beta\right],$
where $C$ is an absolute constant.
@@
$\frac{1}{n}|\pi_{k,f}(z_{i})-\pi_{k-1,f}(z_{i})|$ $\leq\frac{1}{n}|\pi_{k,f}(z_{i})-\pi_{k-1,f}(z_{i})|+\frac{1}{n}|f(z_{i})-f_{k-1}(z_{i})|$
$\leq\frac{\varepsilon_{k}+\varepsilon_{k-1}}{n}$
$\leq\frac{3\varepsilon_{k}}{n}.$
Hence, $\frac{1}{n}\sum_{i\in[n]}x_{i}[\pi_{k,f}(z_{i})-\pi_{k-1,f}(z_{i})]$ is a $(3\varepsilon_{k}/\sqrt{n})$-subGaussian random variable. We conclude that
$\mathbb{E}\left[\sup_{f\in\cal F}\frac{1}{n}\bigg|\sum_{i\in[n]}x_{i}[\pi_{k,f}(z_{i})-\pi_{k-1,f}(z_{i})]\bigg|\right]\leq C\frac{\varepsilon_{k}\sqrt{\log(\mathcal{N}(\mathcal{F},\|\cdot\|_{\infty};\varepsilon_{k}))}}{\sqrt{n}},$
@@
$\leq C^{\prime}\inf_{\alpha\geq 0}\left[\alpha+\frac{1}{\sqrt{n}}\int_{\alpha}^{\infty}\sqrt{\log(\mathcal{N}(\mathcal{F},\|\cdot\|_{\infty};\beta))}d\beta\right].$
@@
From [gottlieb2016adaptive] (Lemma 4.2), we know that
@@
where $\mathcal{F}_{L}$ is the class of $L$-Lipschitz functions mapping $\Omega$ to [1,1][-1,1]. It is easy to see that $D\mathcal{F}_{1/D}={\cal F}_{BL}$, since ${\cal F}_{BL}$ contains all the $1$-Lipschitz functions mapping $\Omega$ to [D,D][-D,D]. Then, for any $\beta\in[0,D]$, $\log(\mathcal{N}({\cal F}_{BL},\|\cdot\|_{\infty};\beta))$ can be bounded as follows
@@
where the last inequality uses the fact that $\beta\in[0,D]$. Replacing this expression in the bound given by Lemma E.4 we get
$G_{n}({\cal F}_{BL})$ $\leq\inf_{\alpha\geq 0}\left(\alpha+\frac{1}{\sqrt{n}}\int_{\alpha}^{D}\left(\frac{5D}{\beta}\right)^{d/2}\sqrt{\log\left(\frac{8D}{\beta}\right)}d\beta\right)$
$\leq\inf_{\alpha\geq 0}\left(\alpha+\frac{\sqrt{\log(8D/\alpha)}}{\sqrt{n}}\int_{\alpha}^{D}\left(\frac{5D}{\beta}\right)^{d/2}d\beta\right)$
$=\inf_{\alpha\geq 0}\left(\alpha+\frac{\sqrt{\log(8D/\alpha)}(5D)^{d/2}}{\sqrt{n}}\int_{\alpha}^{D}\beta^{-d/2}d\beta\right).$
@@
    $\int_{\alpha}^{D}\beta^{-d/2}d\beta=2(\sqrt{D}-\sqrt{\alpha})\leq 2\sqrt{D}.$
@@
    $\int_{\alpha}^{D}\beta^{-d/2}d\beta=\log(D)-\log(\alpha)=\log(D/\alpha)\leq\log(8D/\alpha).$
@@
    $\int_{\alpha}^{D}\beta^{-d/2}d\beta=\frac{\alpha^{1-d/2}-D^{1-d/2}}{d/2-1}\leq\frac{\alpha^{1-d/2}}{d/2-1}.$
@@
The convergence of empirical measures in $1$-Wasserstein is a topic of interest on its own, that has been studied in the past. We obtain the following corollary from [lei2020empiricalwasserstein]. We note the results in there hold more generally for Banach Spaces.
Lemma E.5 (Corollary of Theorem 3.1 in [lei2020empiricalwasserstein]).
@@
Appendix F Simulations with histogram from Section 4.2
We follow the same simulation setup from Section 6. Namely, we consider the sample space to be the unit $\ell_{2}$ ball in $\mathbb{R}^{2}$. We use privacy parameters $\varepsilon=1,\delta=10^{-4}$. We set $T=2\log(n\varepsilon)$, and the noise $\sigma$ is computed with the analytic Gaussian mechanism ([balle2018improvingGaussianmech], Theorem 8). The rest of the parameters are set as indicated in Theorem 4.1. We run Algorithm 2, with the only difference that we post-process noisy histograms truncating at $H=0$ and then re-normalize. In Figure 7, we refer to this algorithm as ‘PE’.
@@
Figure 7: Comparison of the performance of PE with different histograms. Both algorithms, ‘PE’ and ‘PE with Laplace noise+thresholding’ are described in Section F. Left: the sensitive dataset is distributed uniformly in an $\ell_{2}$ ball of radious $0.02$ around the origin (heavily clustered), leading ‘PE with Laplace noise+thresholding’ to preserve the votes signal in the noisy NN histograms, outperforming ‘PE’. Right: the sensitive dataset is distributed uniformly in an $\ell_{2}$ ball of radious $0.5$ around the origin (not clustered). The votes signal in the noisy NN histograms is lost by ‘PE with Laplace noise+thresholding’ during the thresholding step. Hence, it performs poorly. Results are averaged over 100 runs
Appendix G Extension to Banach Spaces
@@
Theorem G.1 (Convergence of PE).
Let $(\Omega,\rho)$ be a (sample) Banach space. Suppose $\Omega$ is compact. Let $1>\gamma>0$, $v\in\mathbb{N}$ and $\alpha>0$. $S_{T}$ be the output of Algorithm 2 run on input $S\in\Omega^{n},T\in\mathbb{N},n_{s}\in\mathbb{N},\sigma>0,\rho(\cdot,\cdot)$. If $\operatorname{Variation\_API}(\cdot)$ is an $(\gamma,v,\alpha)$-API for $S$, then
@@
The upper bound on the Gaussian complexity that we provided in Lemma E.4 can be used to find $\hat{G}_{vn_{s}}({\cal F}_{BL})$, since it works for general metric spaces, and the upper bound $\hat{W}_{1}(n_{s}))$, can be found in [lei2020empiricalwasserstein].
The proof of Theorem G.1 follows similarly to Theorem 4.1. We provide it below for completeness.
Proof of Theorem G.1.
Note that for any t=0,,T1t=0,...,T-1
@@
where the inequality follows from $D_{BL}(\hat{\mu}_{t+1},\tilde{\mu}_{t+1})\geq D_{BL}(\tilde{\mu}_{t+1},\mu^{\prime}_{t+1})$ by definition of $\mu^{\prime}_{t+1}$. Denote by $\mathbb{E}_{t}[\cdot]$ the expectation when conditioning on the randomness up to iteration $t$ of PE. Since $\operatorname{Variation\_API}(\cdot)$ is an $(\gamma,v,\alpha)$-API, then Lemma E.2 gives
@@
Further, since $|V_{t}|=vn_{s}$, Lemma E.3 and the definition of $\hat{G}_{n_{s}}({\cal F}_{BL})$ imply
@@
Appendix H Missing proofs
@@
  • If $\hat{\mu}[i]>0$, then $\tilde{\mu}[i]=\hat{\mu}[i]+L_{i}$, where $L_{i}\sim\text{Lap}(2/n\varepsilon)$. If $\tilde{\mu}[i]<2\log(1/\delta)/(n\varepsilon)+1/n$ (i.e the coordinate is small), then it is truncated: $\tilde{\mu}[i]=0$ .
@@
[…]
@@
For simplicity denote the packing number $\mathcal{M}(\Omega,\rho;2\eta)$ by $M$ in this proof. Consider a $2\eta$-packing of $\Omega$, $\{z_{1},...,z_{M}\}$. Let $S=\{z_{1}\}^{n}$. For $k\in\mathbb{Z}_{+}$, define
@@
It is easy to see that if $S=_{\eta}\mathcal{A}(S)\iff\mathcal{A}(S)\in B(1)$, and if $\mathcal{A}(S)\in B(k)$ for some $k\geq 2$, then $S$ and $\mathcal{A}(S)$ can not be $\eta$-close. Since $\mathbb{P}_{\mathcal{A}}\bigg[S=_{\eta}\mathcal{A}(S)\bigg]\geq 1-\tau$, then
$\mathbb{P}_{\mathcal{A}}\bigg[\mathcal{A}(S)\in\cup_{k\in\mathbb{Z}\cap[2,M]}B(k)\bigg]\leq\tau.$
Furthermore, $B(i)\cap B(j)=\emptyset$ for all $i\neq j\in\mathbb{Z}\cap[2,M]$, so
$\mathbb{P}_{\mathcal{A}}\bigg[\mathcal{A}(S)\in\cup_{k\in\mathbb{Z}\cap[2,M]}B(k)\bigg]=\sum_{k\in\mathbb{Z}\cap[2,M]}\mathbb{P}_{\mathcal{A}}\bigg[\mathcal{A}(S)\in B(k)\bigg].$
Hence, there exists a set $B(k^{*})$ such with $\mathbb{P}_{\mathcal{A}}\bigg[\mathcal{A}(S)\in B(k^{*})\bigg]\leq\tau/(M-1)$. Finally, construct $S^{\prime}=\{z_{1}\}^{n-1}\cup\{z_{k^{*}}\}$. Note that $S$ and $S^{\prime}$ are neighboring datasets. Then, by the inequalities that we have stated and $(\varepsilon,\delta)$-DP, it follows that
$1-\tau\leq\mathbb{P}_{\mathcal{A}}\bigg[S^{\prime}=_{\eta}\mathcal{A}(S^{\prime})\bigg]=\mathbb{P}_{\mathcal{A}}\bigg[\mathcal{A}(S^{\prime})\in B(k^{*})\bigg]\leq e^{\varepsilon}\mathbb{P}_{\mathcal{A}}\bigg[\mathcal{A}(S)\in B(k^{*})\bigg]+\delta\leq\frac{e^{\varepsilon}\tau}{M-1}+\delta.$
Solving for $\varepsilon$, we get $\varepsilon\geq\log((M-1)(1-\tau-\delta)/\tau)$. Finally, note that
$(M-1)(1-\tau-\delta)/\tau\geq M(1-\tau-\delta)/[2\tau]\geq M,$
where the last inequality follows from the fact that $1-\tau-\delta\geq 2\tau$, which is a consequence of our assumption that $\delta<1-3\tau$. This concludes the proof. ∎
@@
Appendix I Details on the Private Signed Measure Mechanism
The Private Signed Measure Mechanism is an algorithm for differentially private synthetic data generation presented in [Vershynin23DPsyntheticdata]. It is designed to work under pure differential privacy. For completeness, we provide an extension of it that works under approximate DP. The algorithm and its proof closely resemble the one from [Vershynin23DPsyntheticdata], with the only difference being that we need to deal with Gaussian noises instead of Laplace in order to achieve approximate DP. As we mentioned in Section 5, PE can be seen as a sequential version of PSMM, and the gaussian mechanisms has better composition guarantees than the Laplace mechanism. Hence, even though changing Gaussian noise by Laplace in PSMM does not make a big difference, it does make a more notorious difference for PE when seen as a sequential version of PSMM.
@@ Appendix I Details on the Private Signed Measure Mechanism
[…]
@@
Theorem I.1 (Guarantees of PSMM under approximate DP).
@@
Remark I.1.
When running PSMM, we can select $n_{s}$ arbitrarily large, by the post-processing property of DP. Hence, we can safely assume that the error term $D\left(\frac{\sqrt{\log(1/\delta)}}{n\varepsilon}\right)^{1/\max\{2,d\}}$ dominates in the utility bound presented in Theorem I.1. Note that, up to logarithmic factors, this bound improves over the bound that we provide for PE in Theorem 4.1 by a factor of $d$. However, it requires $\{\Omega_{i}\}_{i\in[m]}$ to be such that $\max_{i\in[m]}\operatorname{diam}(\Omega_{i})=O(Dm^{-1/\max\{d,2\}})$. Constructing such partition in practice is equivalent to finding an $O(Dm^{-1/\max\{d,2\}})$-net of points $\{\omega_{1},....,\omega_{m}\}$ such that the Voronoi partition induced by them is $\{\Omega_{i}\}_{i\in[m]}$. We explained in Section 5 why such net can be difficult to construct in practice.
@@
[…]
@@
[…]
@@
For the second claim, recall from Corollary E.1 that
@@
Finally, since $S^{\prime}\sim\mu^{n_{s}}$, Lemma E.5 gives the following upper bound on the convergence of the empirical measure in $W_{1}$
@@
Appendix J Impact Statement